OpenStudy (anonymous):

Can someone explain to me Natural Logarithm Functions and help me do this example? Write this equation in exponential form 8=2.08

OpenStudy (whpalmer4):

I'm assuming you know about logarithms in general, and just need to know about the natural logarithm. Speak up if that isn't the case! The natural logarithm is just a logarithm computed to the base \(e \approx 2.718281828\) and is typically written as \(\ln\). By the definition of a natural logarithm, if \(e^{n} = x\) then \(\ln x = n\) by the property that \(\log x^a = a \log x\) and \(\ln e = \ln e^1 = 1\).